OpenStudy (anonymous):

Can someone help me with this, and show me how to solve it? Which function has x-intercepts of 1 and 1 and contains the point (2, 3)?

OpenStudy (anonymous):

Oh and the options are A. y = x^2+3 B. y=x^2-1 C. y=x^2-3 D. y=x^2+1